Calculate the angular momentum of the following systems. Refer to Worksheet 19 for the

moment of inertia of highly-symmetric objects. For all items, you need to convert from

rpm to rad/s first.


Blu-ray disc (inner diameter 0.015 m, outer diameter 0.120 m, mass 0.016 kg) 

spinning at a rate of 10,000 rpm



1
Expert's answer
2021-02-10T18:23:47-0500
"\omega=\frac{2\pi}{60}10000=1047\frac{rad}{s}"


"L=I\omega=0.5m\omega(b^2-a^2)\\
L=0.5(0.016)(1047)(0.12^2-0.015^2)=0.12\frac{kg\cdot m^2}{s}"